A recent survey show 60% of all U. S. house hold has wireless phone only. In a random sample of 20 households what is the probability that exactly 10 has wireless only.
Answer #1

Solution

Given that ,

p = 0.60

q = 1 - p = 1 - 0.60= 0.40

n = 20

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 10) = (20 C 10) * 0.6010 * (0.40)20- 10

probability=0.1171